OpenStudy (kewlgeek555):

M-M-MEDAL: Just want you to check some questions to see if I got it right and maybe answer some other questions.

OpenStudy (kewlgeek555):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = one thirdx - 3. y = 3x - 14 y = one thirdx + 2 y = -3x + 22 y = -one thirdx + 8

OpenStudy (kewlgeek555):

It MUST be B, right?

OpenStudy (doggy):

It is B

OpenStudy (kewlgeek555):

Thans, can you check another one for me, please? Thanks.

OpenStudy (doggy):

I'll try

OpenStudy (kewlgeek555):

Okay. Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 20? The graph of y = f(x) will shift left 20 units. The graph of y = f(x) will shift down 20 units. The graph of y = f(x) will shift up 20 units. The graph of y = f(x) will shift right 20 units I chose C because I know that the y-intercept is the vertical part of the coordinate plane and positive is up.

OpenStudy (doggy):

I don't know how to do this one, sorry :[

OpenStudy (kewlgeek555):

It's okay. "-D

OpenStudy (kewlgeek555):

Anyone else?

OpenStudy (kewlgeek555):

@atlas @mathstudent55

OpenStudy (atlas):

since the line is parallel to y =(1/3)x -3 ; the slope of this line will remain the same you can write the new line as y = (1/3)x +c

OpenStudy (atlas):

You know at least one point that this line passes through So if you put the value of x and y from that point ->you can get c

OpenStudy (kewlgeek555):

Are we talking about C?

OpenStudy (kewlgeek555):

C. Pfft, sorry. I meant the second question.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

no a

OpenStudy (atlas):

C is right in my opinion

OpenStudy (kewlgeek555):

Yeah, I think its C, too because it is the y-intercept...

OpenStudy (atlas):

You can do a simple test to find out..........let f(x) be a constant lets say 1 and see how y changes

OpenStudy (kewlgeek555):

Okay, I have this question that I have been looking at for too long: The following function defines a recursive sequence. f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (atlas):

Just keep on putting the value of n=1,2,3 and so on

OpenStudy (atlas):

in the function f(n)

OpenStudy (kewlgeek555):

Oh no, never mind. I think I know now...It is *-3, ight? Er, multiplying by -3. So 12 * -3 = -36...No, that is not the answer. GR.

OpenStudy (atlas):

you are leaving the 2nd term of f(n)

OpenStudy (atlas):

f(n) = -3•f(n -1) - 2•f(n - 2)

OpenStudy (atlas):

f(3) =-3*f(2)-2*f(1)

OpenStudy (kewlgeek555):

I think I simplified it to f(3) = -5 * f(2) * f(1)...

OpenStudy (atlas):

nope how can you do that??

OpenStudy (kewlgeek555):

Gr. -covers eyes- I did the like terms thingie. Sigh.

OpenStudy (atlas):

but there isn't any like term here..........

OpenStudy (kewlgeek555):

llllllllllllll. Then I dont now how to do this.